OpenStudy (laddiusmaximus):

Find the equation of the line perpendicular to the line4x+8y=16 and through the midpoint of AB where A(-8,12) and B(6,10) Please go step by step.

OpenStudy (anonymous):

can you fist find the midpoint of AB ?

OpenStudy (anonymous):

*first

OpenStudy (anonymous):

do u know the midpoint formula?

OpenStudy (laddiusmaximus):

using the midpoint formula, yes (-1, 11)

OpenStudy (anonymous):

ok... good... now that we have a point, we need the slope of the line we need...

OpenStudy (anonymous):

it needs to be PERPENDICULAR to the line 4x + 8y = 16. can you find the slope of this line?

OpenStudy (laddiusmaximus):

should be -1/2

OpenStudy (anonymous):

yes.... that's the slope of 4x + 8y = 16..... now the line needs to be perpendicular to a slope of -1/2.. what slope would that be?

OpenStudy (laddiusmaximus):

2/1 or just 2

OpenStudy (anonymous):

corrrect... :)

OpenStudy (anonymous):

so we have a point, (-1, 11) and the slope, m=2 we can put that info in point-slope form: \(\large y-y_1=m(x-x_1) \)

OpenStudy (anonymous):

what form does the line need to be in?

OpenStudy (anonymous):

standard form? (like the given line in the original post?)

OpenStudy (laddiusmaximus):

it is in a different form i think the last answer was x-2y=10 which im not sure what form that is

OpenStudy (anonymous):

ok.... stick the info into point-slope form to find the equation of the line we need. after finding that, we can put that in standard form....

OpenStudy (laddiusmaximus):

ok then its y-11=2(x+1)

OpenStudy (anonymous):

yes.... carry out distributive property on the right side to get: y - 11 = 2x + 2 to get it in standard form bring everything with a variable to the left side and move all constants to the right...

OpenStudy (anonymous):

ie, move the 2x to the left and move the -11 to the right.

OpenStudy (laddiusmaximus):

2x+y=13

OpenStudy (laddiusmaximus):

right?

OpenStudy (laddiusmaximus):

the answer is x+2y=21 have no clue how the book got that

hero (hero):

I'll have to check over the steps

hero (hero):

4x+8y=16 4(x + 2y) = 4(4) x + 2y = 4 2y = 4 - x y = (4 - x)/2 y = 2 - x/2 The slope of the original line is definitely -1/2 so the slope of the perpendicular line should be 2.

hero (hero):

Looks like your book is wrong.

OpenStudy (laddiusmaximus):

wow. and I paid for that book too....figures

OpenStudy (laddiusmaximus):

does it always have to be in standard form? can you run that past me again?

hero (hero):

No, there are other forms. Slope Intercept is another form.